in a scene in an action movie a stut man jumps from the top of one building to the top of another 4.0m away, After a running start, he leps at a velocity of 5.0m/s at an angle 15 with respect to the flat roof. will he make it to he other. roof which is 2.5m shorter then te building he jumps from?
1
Expert's answer
2020-11-05T10:38:47-0500
"h=H-vt\sin{15}-0.5gt^2\\-2.5=5t\sin{15}-0.5(9.8)t^2\\t=0.858\ s"

"x=(5)(0.858)\cos{15}=4.14\ m"

It is greater than 4.0 m.

Answer: yes.